answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Function of an anastomosis?

it is important especially in joints where joint movements can impede major arterial channels to supply blood to the joint, therefore anastomosis helps to insure blood flow. Besides this, anastomosis is important when a major artery has been blocked and the blood then can be carried through anastomotic arteries.

Odds of being killed by toilet?

Negligible, although it can depend on what you do with a toilet.

Of the approx 7 billion people on the planet at present, I would be surprised if a thousand were killed by a toilet. Even a million fatalities would give a probability of less than 0.00015.

One card is drawn from a standard pack of 52 what is the probability that it is either a king or a queen?

P(Card is King) = 4/52

P(Card is Queen) = 4/52

Since these are mutually exclusive events.i.e. they can't happen in a single trial these two probabilities will be added.

P(King or Queen) = 4/52 + 4/52 = 8/52

What is the probabilty of picking an ace in a deck of 52 cards?

If you pick enough cards, the probability is 1. For a single card, picked at random, the probability is 1/13.

Probability is the chance that something will or will not happen?

Usually, probability is the chance that something will happen; but it can be will not happen as well.

How many ways is there to seat six guest at a table for 6?

The first person has a choice of 6 seats, the second one has only 5 choices, the third one 4 choices, the fourth one 3 choices, the fifth one 2 choices a, and the sixth person has to take the one seat that is left.

So the answer is 6! or 6x5x4x3x2x1 = 720 ways.

Number of spades on a ace of spades?

There are three spades on the standard ace of spades. One in each corner and a large one in the middle.

How do you answer 'When something goes wrong who do you tend to assume is at fault' in a job interview?

I'm assuming you mean "on the job" and not "at the interview." Sounds like a trick question. It doesn't matter who is at fault (well, unless the problem is something involving something illegal or grounds for dismissal). The real issue is how the problem is solved. Placing blame solves nothing. Solving the problem (or helping to)shows intiative, that you care about the company and that you're a doer and not just a thinker.

What distance is longer 1 mile 1000 meters or 10000 centimeters?

1 mile = 1 mile

1000 meters = 0.6213711922373 miles

10000 centimeters = 0.06213711922373 miles

1 mile is the longest

1000 meters is 0.6213712 of a mile

1 mile = 1.609 km

1km = 1000m

100m=10000cm

therefore 1 mile is the longest

What is an example of equality or theoretical justice?

An example of equality or theoretical justice is the concept of equal access to education for all individuals, regardless of their socioeconomic background, race, or gender. This principle asserts that everyone should have the same opportunities to gain knowledge and skills, which promotes fairness and enables individuals to improve their circumstances. In a just society, policies would be implemented to eliminate barriers and ensure that all students receive the resources and support they need to succeed.

How many ways can you arrange 6 things?

This is a simple math permutation. the answer is 6 factorial which means 6 * 5 * 4 * 3 * 2 * 1 (aka 720)

Small sample size?

Placing a question mark at the end of a phrase does not make it a sensible question. Try to use a whole sentence to describe what it is that you want answered.

If something happens at your work place and they close down for a week do you still get paid?

You will need to apply for unemployment that week of shut down and wait several weeks to get paid.

How many ways can you arrange the letters in the word rainbow?

Since there are 7 letters in the word rainbow, and all are different, the answer is 7! or 5040.

Trending Questions
When two dice are rolled what is the probability that both will not be the same number? What is the probability of not choosing a green marble from a jar containing 5 red 6 green and 4 blue marbles? If a couple wants to have two children what is the probability that the first will be a girl and the second will be a boy? What is the probability of choosing a red marble or a green marble out of a bag containing 5 blue marbles 3 red marbles and 1 green marble? A jar contains 8 red marbles 9 blue marbles and 6 green marbles Two marbles are picked at random What is the probability that one is green and one is blue? What is the difference between arterial and venal blood? If a bag contains 20 chips labeled with the numbers 1 through 20 and one chip is selected at random. what is the probability that the chip has the number 17 on it? Give an example of how standard deviation can be useful Also why is underestimating the standard deviation as in the case with the Range Rule Thumb a better method than overestimating? If you flip four quarters all at once what is the likelihood that you will get 2 heads and 2 tails? How many 2 letter arrangements are possible from 26 letters if no letter can be repeated? Throw a dice and toss a coin what is the probability that a number less than 3 and a tail result? Ace combat 4 how do you eject? What is the probability of getting four heads when flipping four identical coins at once? Sample space of tossing a die once? How many red spades in a deck of cards? What is the probability of rolling an even number if you will roll a six-sided number cube with sides numbered 1 through 6? What is the Probability density function of Poisson distribution? What is the probability that you will select three odd-numbered cards in a row from a bag of 7 cards? What are the criteria for selecting between bidders? Features on the probability scale?